# SparseVLM: Visual Token Sparsification for Efficient Vision-Language Model Inference

### v3

@@ 5.4 Qualitative Visualization
As shown in Figure 6, we visualize SparseVLM on various VQA questions. From left to right, we visualize the results after we apply token pruning to different layers. As the number of layers increases, more tokens are pruned and the Region of Interest (ROI) is gradually refined. The model systematically reduces less relevant image information while retaining key tokens closely tied to the question. The visualization reveals that SparseVLM, although discarding some overall image details, effectively retains essential visual tokens. These preserved tokens encapsulate the features necessary for answering the question, focusing on more relevant visual regions through their interaction with the question. More cases are in the Appendix A.7.
@@ 6 Conclusion
This paper introduced a text-aware training-free token optimization approach called SparseVLM which significantly decreased the test-time computations of various VLMs. Unlike prior methods, SparseVLM optimized VLMs without introducing extra parameters and fine-tuning costs. We achieved a more compact visual representation by employing the rank of attention matrices to determine pruning ratios and by recycling the pruned tokens via the reconstruction mechanism to reduce the information loss. Experiments demonstrated that e.g. the LLaVA when equipped with SparseVLM achieved 37.0% reduction in latency with a compression ratio of 77.8% while maintaining 97% of the original accuracy. Moreover, our method exceeded FastV accuracy by 34.4% in video understanding tasks. Our SparseVLM can provide practical benefits for deploying off-the-shelf VLMs on edge devices and in the cloud setting.
@@
Our SparseVLM offers tangible advantages for implementing off-the-shelf VLMs on edge devices and in cloud environments. While our work carries no discernible societal implications, we see no necessity to underscore this aspect in the present context.
Appendix A Appendix
Figure 7: Comparison of visual redundancy in different vision tasks.
A.1 The Redundancy of Visual Tokens in VLM
In non-textual tasks like classification or segmentation, downsampling is commonly employed to decrease visual redundancy and enhance model training efficiency. Figure 7 illustrates this process, showing the reduction of tokens from 1166 to 576 in a downsampled image, resulting in a $50\%$ efficiency boost but a $15\%$ information loss (entropy decreased from 7.44 to 6.13). This trade-off is acceptable for such tasks. Conversely, for text-related tasks like visual question answering (VQA), which involve text and vision modalities, a distinct approach is needed. Highlighting the most information-dense text ($88\%$ of total text) alongside the region pertinent to the query in the image ($38\%$ of total image), we observe that image information is typically sparser than textual data. Hence, our SparseVLM method incrementally prunes visual token redundancy, maintaining crucial information for task accuracy. This strategy enhances model efficiency and effectiveness.
A.2 Compatibility with FlashAttention
@@
A.3 Computing Budget Detailed Estimation
@@ A.4 Dataset
GQA. (Hudson & Manning, 2019) The GQA benchmark is composed of three parts: scene graphs, questions, and images. The image part contains images, as well as the spatial features of images and the features of all objects in images. The questions in GQA are designed to test the understanding of visual scenes and the ability to reason about different aspects of an image.
MMBench. (Liu et al., 2023b) The MMBench benchmark comprehensively evaluates the model’s overall performance across multiple dimensions. It includes three levels of ability dimensions. The first level (L-1) consists of two main abilities, perception and reasoning. The second level (L-2) expands based on the first level, including six sub-abilities. The third level (L-3) further refines the second level, encompassing 20 specific ability dimensions. This hierarchical structure enables a granular and comprehensive evaluation of the model’s various capabilities.
@@
MMVet. (Yu et al., 2023) The MMVet benchmark is designed based on the insight that the intriguing ability to solve complicated tasks is often achieved by a generalist model being able to integrate different core vision-language capabilities. MM-Vet defines 6 core VL capabilities and examines the 16 integrations of interest derived from the capability combination.
@@
A.5 Implementation Details
All of our experiments are conducted on a single NVIDIA A100-80G GPU. The implementation is carried out in Python 3.10, utilizing PyTorch 2.1.2, CUDA 11.8, and transformers 4.31.0. The inference follows the evaluation settings established by LLaVA (Liu et al., 2024). For LLaVA-1.5-7/13B, Mini-Gemini (MGM), and Qwen-VL, we follow the same inference setting as the original paper as it is publicly available111github.com/haotian-liu/LLaVA 222github.com/dvlab-research/MGM  [^3] https://github.com/QwenLM/Qwen-VL. For video understanding tasks, we adopt the same inference setup as the original Video-LLaVA code base444github.com/PKU-YuanGroup/Video-LLaVA., as it is publicly available.
A.6 Efficiency Details
We present a comparative efficiency analysis of SparseVLM, the baseline, and FastV (Chen et al., 2024) during the inference phase in Table 1. In this section, we provide additional details on the CUDA time during the inference phase. Following VoCo-LLaMA (Ye et al., 2024), we primarily consider the following components that contribute to the reported CUDA time: image encoding time (if applicable), KV cache load time (if applicable), and transformers forward time. We exclude other computational times that are not dependent on the model itself and the caching strategy, such as model loading time, from the CUDA time measurement. Specifically, the attention operation is implemented by FlashAttention (Dao et al., 2022).
A.7 More Sparsification Visualization
Figure 8 showcases a diverse array of visualization examples that demonstrate the application of SparseVLM across a spectrum of Visual Question Answering (VQA) prompts. These visualizations offer a deeper insight into how SparseVLM processes and responds to different types of queries posed in a visual context.
Figure 8: More visualization examples of SparseVLM on different VQA prompts.
A.8 More Detailed Efficiency Analysis
@@
Figure 9: Trade-offs for SparseVLM on LLaVA: (a) Latency vs. Accuracy, and (b) FLOPs vs. Accuracy. Both show comparisons of random sparse, SparseVLM, and baseline.
Figure 10: Trade-offs for SparseVLM on MGM: (a) Latency vs. Accuracy, and (b) FLOPs vs. Accuracy. Both show comparisons of random sparse, SparseVLM, and baseline.
Figure 11: Trade-offs for SparseVLM on Video-LLaVA: (a) Latency vs. Accuracy, and (b) Token budget vs. Accuracy. Both show comparisons of SparseVLM, FastV, and baseline.

### v4

@@ 5.4 Qualitative Visualization
As shown in Figure 6, we visualize SparseVLM on various VQA questions. From left to right, we visualize the results after we apply token pruning to different layers. As the number of layers increases, more tokens are pruned and the Region of Interest (ROI) is gradually refined. The model systematically reduces less relevant image information while retaining key tokens closely tied to the question. The visualization reveals that SparseVLM, although discarding some overall image details, effectively retains essential visual tokens. These preserved tokens encapsulate the features necessary for answering the question, focusing on more relevant visual regions through their interaction with the question. More cases are in the Appendix H.
@@ 6 Conclusion
This paper introduced a text-aware training-free token optimization approach called SparseVLM which significantly decreased the test-time computations of various VLMs. Unlike prior methods, SparseVLM optimized VLMs without introducing extra parameters and fine-tuning costs. We achieved a more compact visual representation by employing the rank of attention matrices to determine pruning ratios and by recycling the pruned tokens via the reconstruction mechanism to reduce the information loss. Experiments demonstrated that e.g. the LLaVA when equipped with SparseVLM achieved 37.0% reduction in latency with a compression ratio of 77.8% while maintaining 97% of the original accuracy. Moreover, our method exceeded FastV accuracy by 14.7% in video understanding tasks. Our SparseVLM can provide practical benefits for deploying off-the-shelf VLMs on edge devices and in the cloud setting.
Acknowledgments
This work was supported by the National Science and Technology Major Project (No. 2022ZD0117800) and by the National Natural Science Foundation of China under Grant 62472008.
@@ Acknowledgments
Our SparseVLM provides practical advantages for deploying off-the-shelf large vision-language models on edge devices and cloud platforms. While our work does not present any evident societal implications, we believe it is unnecessary to emphasize this aspect in the current context.
Appendix
Appendix A The Redundancy of Visual Tokens in VLMs
In non-textual tasks, such as classification or detection, downsampling is commonly used to reduce visual redundancy and enhance model training efficiency (Zhang et al., 2024b). Figure 7 illustrates this process, showing the reduction of tokens from 1166 to 576 in a downsampled image, resulting in a $50\%$ efficiency boost but a $15\%$ information loss (entropy decreased from 7.44 to 6.13). This trade-off is acceptable for such tasks. Conversely, for text-related tasks like visual question answering (VQA), which involve both text and vision modalities, a distinct approach is required. Highlighting the most information-dense text ($88\%$ of total text) alongside the region pertinent to the query in the image ($38\%$ of total image), we observe that image information is typically sparser than textual data. Hence, our SparseVLM method incrementally prunes visual token redundancy, maintaining crucial information for task accuracy. This strategy enhances model efficiency.
Figure 7: Analysis of visual redundancy in different vision tasks.
Appendix B Compatibility with FlashAttention
@@
Appendix C Computing Budget Detailed Estimation
@@ Appendix D Dataset
GQA. (Hudson & Manning, 2019) The GQA is composed of three parts: scene graphs, questions, and images. The image part contains images, as well as the spatial features of images and the features of all objects in images. The questions in GQA are designed to test the understanding of visual scenes and the ability to reason about different aspects of an image.
MMBench. (Liu et al., 2024c) The MMBench benchmark comprehensively evaluates the model’s overall performance across multiple dimensions. It includes three levels of ability dimensions. The first level (L-1) consists of two main abilities, perception and reasoning. The second level (L-2) expands based on the first level, including six sub-abilities. The third level (L-3) further refines the second level, encompassing 20 specific ability dimensions. This hierarchical structure enables a granular and comprehensive evaluation of the model’s various capabilities.
@@
MMVet. (Yu et al., 2024) The MMVet benchmark is designed based on the insight that the intriguing ability to solve complicated tasks is often achieved by a generalist model being able to integrate different core vision-language capabilities. MM-Vet defines 6 core VL capabilities and examines the 16 integrations of interest derived from the capability combination.
@@
Appendix E Implementation Details
All of our experiments are conducted on a single NVIDIA A100-80G GPU. The implementation is carried out in Python $3.10$, utilizing PyTorch $2.1.2$, CUDA $11.8$, and transformers $4.31.0$. The inference follows the evaluation settings established by LLaVA (Liu et al., 2024b). For LLaVA-1.5-7/13B, Mini-Gemini (MGM), and Qwen-VL, we follow the same inference setting as the original paper as it is publicly available111github.com/haotian-liu/LLaVA 222github.com/dvlab-research/MGM  [^3] https://github.com/QwenLM/Qwen-VL. For video understanding tasks, we adopt the same inference setup as the original Video-LLaVA code base444github.com/PKU-YuanGroup/Video-LLaVA., as it is publicly available.
Appendix F Efficiency Details
We present a comparative efficiency analysis of SparseVLM, the baseline, and FastV during the inference phase in Table 1. In this section, we provide additional details on the CUDA time during the inference phase. Following VoCo-LLaMA (Ye et al., 2025), we primarily consider the following components that contribute to the reported CUDA time: image encoding time (if applicable), KV cache load time (if applicable), and transformers forward time. We exclude other computational times that are not dependent on the model itself and the caching strategy, such as model loading time, from the CUDA time measurement. Specifically, the attention operation is implemented by FlashAttention (Dao et al., 2022).
Appendix G More Detailed Efficiency Analysis
@@
Figure 8: Trade-offs for SparseVLM on LLaVA: (a) Latency vs. Accuracy, and (b) FLOPs vs. Accuracy. Both studies demonstrate comparisons among random sparse, our SparseVLM, and baseline models.
Figure 9: Trade-offs performance for SparseVLM on MGM: (a) Latency vs. Accuracy, and (b) FLOPs vs. Accuracy. Both studies demonstrate comparisons among random sparse, our SparseVLM, and baseline models.
Figure 10: Trade-offs for SparseVLM on Video-LLaVA: (a) Latency vs. Accuracy, and (b) Token budget vs. Accuracy. Both studies demonstrate comparisons among random sparse, our SparseVLM, and baseline models.
Appendix H More Sparsification Visualization
Figure 11 showcases a diverse array of visualization examples that demonstrate the application of SparseVLM across a spectrum of visual question-answering (VQA) prompts. These visualizations offer a deeper insight into how our SparseVLM processes and responds to different types of queries posed in a visual context.
Figure 11: More visualization examples of SparseVLM on different prompts. Best viewed in color.